OpenStudy (anonymous):

Can someone help please? Solve each of the following equations 7x+7(-2x-3)=-22-6x

OpenStudy (austinl):

You need to distribute, combine like terms, get variables on one side, and constants on the other.

OpenStudy (anonymous):

x=1

OpenStudy (austinl):

@Elieroth We don't just give answers here without proper explanation. For future reference, http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

ok I'm sry

OpenStudy (anonymous):

Thanks Austin

OpenStudy (austinl):

No problem, do you still need help, or do you understand how to get to the answer?

OpenStudy (anonymous):

can you run through step by step please?

OpenStudy (austinl):

Sure. \(7x+7(-2x-3)=-22-6x\) Distribute. \(7x-14x-21=-22-6x\) Then combine like terms. \(-7x-21=-22-6x\) Variables on one side, constants on the other. \(-21-(-22)=-6x-(-7x)=-21+22=-6x+7x\) Therefore! \(1=x\)

OpenStudy (anonymous):

oh I see where I messed up Thanks so much!

OpenStudy (austinl):

No problemo!

OpenStudy (anonymous):

Simplifying 7x + 7(-2x + -3) = -22 + -6x Reorder the terms: 7x + 7(-3 + -2x) = -22 + -6x 7x + (-3 * 7 + -2x * 7) = -22 + -6x 7x + (-21 + -14x) = -22 + -6x Reorder the terms: -21 + 7x + -14x = -22 + -6x Combine like terms: 7x + -14x = -7x -21 + -7x = -22 + -6x Solving -21 + -7x = -22 + -6x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '6x' to each side of the equation. -21 + -7x + 6x = -22 + -6x + 6x Combine like terms: -7x + 6x = -1x -21 + -1x = -22 + -6x + 6x Combine like terms: -6x + 6x = 0 -21 + -1x = -22 + 0 -21 + -1x = -22 Add '21' to each side of the equation. -21 + 21 + -1x = -22 + 21 Combine like terms: -21 + 21 = 0 0 + -1x = -22 + 21 -1x = -22 + 21 Combine like terms: -22 + 21 = -1 -1x = -1 Divide each side by '-1'. x = 1 Simplifying x = 1

OpenStudy (anonymous):

Thanks Elie :]

OpenStudy (anonymous):

u r welcome